The expected number of tickets a student is going to have is 15 within their lifetime with a standard deviation of 6.2. Would having 0 tickets be​ unusual? Explain using the Range Rule of Thumb. Round to the nearest whole number. What is the minimum number of tickets students will​ get? nothing What is the maximum number of tickets students will​ get? nothing Would it be unusual for a student to have 0​ tickets?

A. ​Yes, 0 is outside the interval.

B. ​No, 0 is within the interval.

C. ​Yes, 0 is within the interval.

D. ​No, 0 is outside the interval.
Answer #1

Answer)

According to the range of thumb rule

Upper limit = mean + 2*sd = 15 + 2*6.2 = 27.4

Lower limit = mean - 2*sd = 15 - 2*6.2 = 2.6

Anything above 27.4 or below 2.6 is considered as unusual.

So, as 0 is less than 2.6, it is not usual.

So the correct answer is

Yes, 0 is outside the interval